OpenStudy (anonymous):

How do I find the difference quotient? I already have the function of x and the equation and the restriction. How do I start?

OpenStudy (raden):

share ur question, please...

OpenStudy (anonymous):

i have already plugged in (3+h) for x. what do i do next? @RadEn

OpenStudy (raden):

about derivative by using limit methods, right ?

OpenStudy (anonymous):

\(f(6)=5\) so we keep that number

OpenStudy (anonymous):

then compute \[f(6+h)=5+6(6+h)-(6+h)^2\] to get \[5+36+6h-(36+12h+h^2)\] \[=5-6h-h^2\] then subtract 5 so that \(f(6+h)-f(6)=-6h-h^2\) divide by \(h\) and get \(-6-h\)

OpenStudy (anonymous):

so i don't divide by H until i have plugged in both 6+h and 6 and simplified both? @satellite73

OpenStudy (raden):

smart way : f(x) = 5+6x-x^2 f'(x) = 6-2x if u want find value of f'(6), u just plug in x=6 into f'(x) so, f'(6)=6-2(6)=6-12= ????????? but if u want using limit's method, like @satellite73 said

OpenStudy (anonymous):

you do not divide by \(h\) until the end one reason is that you cannot divide by \(h\) until you have simplified the numerator so that everything without an \(h\) in it is gone

OpenStudy (anonymous):

Thank you!! both of your answers were very helpful!

OpenStudy (anonymous):

yw